In the vertical jump, an athlete starts from a crouch and jumps upward as high as possible. Even the best athletes spend little more than 1.00 s in the air (their “hang time”). Treat the athlete as a particle and let Ymax be his maximum height above the floor. To explain why he seems to hang in the air, calculate the ratio of the time he is above Ymax/2  to the time it takes him to go from the floor to that height. Ignore air resistance.


1
Expert's answer
2021-02-15T00:55:50-0500

First, we need to get an expression for "v_o". When the athlete reaches the maximum height, v = 0. So,

                                                               "v^2 = v_o^2-2gy_{max}"

                                                                   "0 = v_o^2-2gy_{max}"

                                                                "v_o=\sqrt{2gy_{max}}"

The time the athlete spends above "\large\frac{y_{max}}{2}" :

For this case,  we are interested in the motion from "\large\frac{y_{max}}{2}" up and back to "\large\frac{y_{max}}{2}" again so

"y = y _o=\large\frac{y_{max}}{2}"

                                    "y -y_o = v_0t_{above}- \frac{1}{2}gt^2_{above}"

                                          "0= v_0t_{above}- \frac{1}{2}gt^2_{above}"

                                              "v_o = \frac{1}{2}gt_{above}"

                                                 "\therefore t_{above} = \large\frac{2v_o}{g}"

Substituting for "v_o" we get:

                                                       "t_{above} = \large\frac{\sqrt{2gy_{max}}}{g}"

The time the athlete takes to reach "\large\frac{y_{max}}{2}" jumping from the ground:

                                                    "y -y_o = v_0t_{below}- \frac{1}{2}gt^2_{below}"

                                                  "\frac{y_{max}}{2} = v_0t_{below}- \frac{1}{2}gt^2_{below}"

Rearranging the equation and multiplying by 2 to put the equation into the quadratic form:

                                            "gt^2_{below}-2v_ot_{below}+y_{max} = 0"

Now we have a quadratic equation with "a = g, \space b = -2v_o \space and \space c = y_{max}"

                                    "t_{below} = \large\frac{-b\pm \sqrt{b^2-4ac}}{2a}"

                                   "t_{below} = \large\frac{-2v_o\pm \sqrt{4v_o^2-4gy_{max}}}{2g}"

                                        "t_{below} = \large\frac{-v_o\pm \sqrt{v_o^2-gy_{max}}}{g}"

Substituting "v_o = \sqrt{2gy_{max}}" we get:

                                          "t_{below} = \large\frac{\sqrt{2gy_{max}}\pm \sqrt{2gy_{max}-gy_{max}}}{g}"

                                         "t_{below} = \large\frac{\sqrt{2gy_{max}}\pm \sqrt{gy_{max}}}{g}"

  We take the value with minus sign in middle of the numerator. The athlete pass by the point

"\large\frac{y_{max}}{2}" twice, once when rising up and again when falling down the minus for the rising up event and the plus for the falling down event.

                                  "\therefore" "t_{below} = \large\frac{\sqrt{2gy_{max}}\pm \sqrt{gy_{max}}}{g} = \frac{\sqrt{gy_{max}}}{g}(\sqrt{2}-1)"

     So, The ratio is:

       "\large\frac{t_{above}}{t_{below}}" "= \large\frac{2\sqrt{2gy_{max}}}{g} * \frac{g}{\sqrt{gy_{max}}(\sqrt2 - 1)}"  "= \large\frac{2\sqrt2}{\sqrt2-1}" "= 6.83"

The athlete spends more time above "\large\frac{y_{max}}{2}" more than below it by a factor of 13.66, so they seem to be hanging in the air